CC Calculate g(x), where g(2) is the inverse of f(x) = 2+1 g(x) =
Answer #1

f(x) = х atl let y x X + Interchange X and Y X = Y Y+1 x(4+1) = y XY+X = Y (x-1) Y = -* Y = X 1-X & f(x) X 1-X Sk) = Х 1-X S